OpenStudy (anonymous):

Express your answer in lowest terms 4/5 x 3/8

OpenStudy (anonymous):

no give the answer

OpenStudy (phi):

\[ \frac{4}{5}\cdot \frac{3}{8} = \frac{4 \cdot 3}{5\cdot 8} \]

OpenStudy (phi):

notice we can divide the top by 4 and we get 1*3 up top and we can divide the bottom by 4 , because 8/4 is 2, and the bottom is 5*2 now simplify \[ \frac{1\cdot 3}{5 \cdot 2} \]

OpenStudy (anonymous):

answer is 3/10

OpenStudy (anonymous):

give medall